How can ionic product be determined using conductivity measurements? Explain.
1
Expert's answer
2020-01-20T05:11:18-0500

"\Lambda_m=\Lambda^0_m -K" "(c)^{1\/2}"

where "{\displaystyle \Lambda _{m}^{0}}" is known as the limiting molar conductivity, K is an empirical constant and c is the electrolyte concentration. (Limiting here means "at the limit of the infinite dilution".)


so "\alpha" (dissociation constant) can be calculated by formula "\Lambda" "_m/\Lambda_m^0"

let's us take a example :

AB(s) "\leftrightarrow" A+(aq) + B+ (aq)

C 0 0 ......initial

C - C"\alpha" C"\alpha" C"\alpha" .......final


so, Solubility product = [A+][B+] = C2*"\alpha^2"

where C is the concentration and "\alpha" is the dissociation constant which can be calculated from the above formula.